in 1895, the first putting green championship was held. the winner's prize money was $220. in 2016, the winner's check was $1,420,000.

a. what was the percentage increase per year in the winner's check over this period? (2 decimal places rounded)

b. if the winner's prize increases at the same rate, what will it be in 2033? (2 decimal places rounded)

please when showing the answer show what you put on the BA II calculator to get the answers. For a, i got 7.52% which is right, but I'm getting b wrong (i'm getting 4,870,860.50) so please help
Answer #1

(a) The question can be solved on the BA II Plus by following the steps described below:

- Input PV = - $ 220, FV = $ 1420000, N = 121 years (2016 - 1895), PMT = 0

- CMPT - > I/Y

- I/Y ~ 0.07519 or 7.52 %

(b) Assuming that the prize money grows at the rate calculated above, the question can be solved on the BA II Plus by following the steps described below:

- Input PV = - $ 220, I/Y = 7.52 % , N = 138 years (2033 - 1895), PMT = 0

- CMPT - > FV

- FV ~ $ 4870334.79